solve the proportion please help
6/a=18/27

Answer by stanbon(75887)

solve the proportion please help
6/a=18/27
----
Invert both sides to get:
a/6 = 26/18
---
a = 6*(26/18)
---
a = 26/3
---
a = 8 2/3
================
Cheers,
Stan H.